Tags:
create new tag
view all tags

Laser Maintenance

LC noticed that the laser was particularly hot after shift on the 21 / 22. PFN was 88.2F at the end of shift. The Nestlab is programmed to 20C (68F). The higher temperature as a detrimental effect on the laser output and stability. LC tested the laser on the afternoon of the 22. When it was off it was 68.2F on the PFN, 15 minutes after turn it was 83F. Typical end of shift values when the laser has been on for a long time are 78F, indicating that this was unusually high.

  • LC emailed Terunuma san around 17:00 / 22 to ask for the water filters to be changed / water topped up.
  • Engineer straight away and they found Neslab fault.
  • 23rd Neslab broken. KEK engineers replacing with new chiller that was spare but going to be used for AA's laser.
  • Internal cooling water check and also low.
  • Internal cooling water topped up by KEK engineers by taking bottom drawer out of PFN, opening perspex cover and pouring in.
  • 14:00 tests of setup.

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r1 - 23 May 2012 - LaurieNevay

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2024 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback